# Universal Ahlfors–David regularity of Steiner trees

### v1

@@ 4.1 Lower bound
We claim that for the vertices of the hypercube $H=\{-1,1\}^{d}\subset\mathbb{R}^{d}$, a Steiner tree is necessarily full. By Lemma 1, a Steiner tree for $H$ lies in the convex hull of $H$, and for any points $x\in H$, $y,z\in\operatorname{conv}H$ the angle $\angle yxz$ is non-obtuse.
@@
$\mathcal{H}^{1}(\mathcal{S}t\cap B_{1}(o))\geq\frac{2(2^{d}-1)}{\sqrt{3d}}$
for this particular $\mathcal{S}t$.
@@ 4.2 Upper bound
By Theorem 3 there exists a set $S$ connecting $\partial B_{r}(x)\cap\mathcal{S}t$ with length at most $8rt_{r}^{\frac{d-2}{d-1}}$, so by Lemma 5 for $X=B_{r}(x)\cap\mathcal{S}t$ and $Y=S$, we have
@@
$t_{r}\geq\frac{L_{r}^{\frac{d-1}{d-2}}}{64}$
for every $r\in[\rho,1)$.
Lemma 6 implies
$L_{R}\geq L_{\rho}+\int_{\rho}^{R}t_{r}\,dr\geq L_{\rho}+\frac{1}{64}\int_{\rho}^{R}L_{r}^{\frac{d-1}{d-2}}dr$ (2)
@@
$f(z)=a+b\int_{\rho}^{z}f(y)^{\alpha}\,dy,\quad a=L_{\rho},\quad b=\frac{1}{64},\quad\alpha=1+\frac{1}{d-2}.$
@@
$L_{\rho}\leq\left(\frac{64(d-2)}{1-\rho}\right)^{d-2}.$
@@
$(q-\rho)t\leq\mathcal{H}^{1}(B_{q}(x)\cap\mathcal{S}t)\leq\left(\frac{64d}{1-q}\right)^{d-2},$
@@
$t\leq\frac{64d^{d-2}(d-1)^{d-1}}{(1-\rho)^{d-1}(d-2)^{d-2}}<\frac{1}{2}\frac{64d^{d-1}}{(1-\rho)^{d-1}}.$
Let $u\in[\rho,q]$ be a radius for which $(B_{u}(x)\cap\mathcal{S}t)=t$. The Steiner set $\mathcal{S}t$ has no terminals inside $B_{u}(x)$, so the number of segments in the forest $\mathcal{S}t\cap B_{u}(x)$ is at most
$2t-3\leq\frac{64d^{d-1}}{(1-\rho)^{d-1}}.$
@@
$\mathcal{H}^{1}(\mathcal{S}t_{\varepsilon}\cap B_{\rho\varepsilon}(x))\leq\mathcal{H}^{1}(\mathcal{S}t\cap B_{\rho\varepsilon}(x))\leq\varepsilon\left(\frac{64d}{1-\rho}\right)^{d-2}.$
@@ 4.3 Planar case
Now consider $\mathcal{S}t_{\varepsilon}=\mathcal{S}t\cap\overline{B_{1-\varepsilon}(o)}$, where $o$ is the origin and $\varepsilon>0$. By Corollary 1 $\mathcal{S}t_{\varepsilon}$ is a finite tree, so formula (1) gives that
$\mathcal{H}^{1}(\mathcal{S}t_{\varepsilon})=\sum_{k=1}^{n}\bar{c_{k}}p_{k},$ (3)
where $\{p_{1},\dots p_{n}\}=\mathcal{S}t\cap\partial B_{1-\varepsilon}(o)$. Lemma 4 states that every summand $\bar{c_{k}}p_{k}$ is $(1-\varepsilon)\cos\alpha_{k}$, where $\alpha_{k}$ is the angle between the radius $op_{k}]$ of $B_{1-\varepsilon}(o)$ and $\mathcal{S}t_{\varepsilon}$ at the point $p_{k}$. Clearly, $\alpha_{k}$ belongs to $0,\pi/2)$, so every summand is positive.
@@
If $x\in\omega$ is an accumulation point of $\mathcal{A}$, then it is an accumulation point from both sides. Formally, for every $\varepsilon>0$ every connected component of $(\omega\cap B_{\varepsilon}(x)\setminus\{x\})$ has a non-empty intersection with $\mathcal{A}$.
@@
The length of a full tree with at least three terminals on a unit circle is at least $\sqrt{3}$.
@@
Suppose the contrary. Then there is a full tree $S$ with the length at most $\sqrt{3}(1-\varepsilon)$; which means that a finite full tree $S_{\varepsilon}$ inside $B_{1-\varepsilon}(o)$ also has length at most $\sqrt{3}(1-\varepsilon)$. We may assume that $S_{\varepsilon}$ has at least 3 terminals, otherwise take smaller $\varepsilon$.
So $S_{\varepsilon}$ is a locally minimal tree and one can apply Lemma 2 (Melzak reduction). By the pigeon-hole principle there are at least two choices of $p_{1},p_{2}$ and $q$, where terminals $p_{1},p_{2}$ that are adjacent to the branching point $q$. Since $S_{\varepsilon}$ is a tree exactly one of the arcs $\breve{p_{1}p_{2}}$ does not contain any other terminals. The interiors of these arcs for two choices are disjoint, so one of them has angular measure smaller than $4\pi/3$, and so the Melzak point $p$ lies outside the circle $\partial B_{1-\varepsilon}(o)$. Also, this arc $\breve{p_{1}p_{2}}$ belongs to the quadrilateral $pp_{1}qp_{2}$ and so Melzak reduction in this case produces a tree $S^{\prime}_{\varepsilon}$. Note that $S^{\prime}_{\varepsilon}\cap B_{1-\varepsilon}(o)$ is a full tree with the length at most $\sqrt{3}(1-\varepsilon)$, so we can repeat the reduction until a tree with 3 terminals $a_{1}$, $a_{2}$ and $a_{3}$. If the diameter of $\{a_{1},a_{2},a_{3}\}$ is at most $\sqrt{3}(1-\varepsilon)$ then they can be covered by an arc $R$ of the angular measure $2\pi/3$; this gives that one of the angles, say $\angle a_{1}a_{2}a_{3}$, is at least $2\pi/3$. This contradicts the existence of a full tree (in this case it is a regular tripod) with terminals $a_{1}$, $a_{2}$ and $a_{3}$. Finally the length of any connected set containing $\{a_{1},a_{2},a_{3}\}$ is at least $\operatorname{diam}\{a_{1},a_{2},a_{3}\}$. ∎
@@ 5 Open questions
We want to understand the behavior of the tight bounds in Theorem 2 and Corollary 2 with respect to $d$ and $1-\rho$.
Also, note that the results by Paolini and Stepanov work in a proper metric space $\mathcal{X}$. The quantitative picture in various metric spaces should be diverse, as we can see even in the planar and non-planar Euclidean cases.
Quantitative regularity of other one-dimensional problems is also of interest. One of the challenging examples is Branched Optimal Transport (Gilbert–Steiner problem), see [1].
@@ Acknowledgments.
The authors are grateful to Ilya Bogdanov for his questions, which led to clarification of the statements. This work started at Summer Scientific Workshop ‘‘Introduction to Modern Mathematics and Related Scientific Projects’’ supported by the Bulgarian Ministry of Education and Science, Scientific Programme ‘‘Enhancing the Research Capacity in Mathematical Sciences (PIKOM)’’, No. D01-88/20.06.2025. The work of P. Prozorov was supported by the Russian Science Foundation (project no. 25-11-00058). Y. Teplitskaya is supported from the French National Research Agency (ANR) under grant ANR-21-CE40-0013-01 (project GeMfaceT).

### v2

@@ 4.1 Lower bound
We claim that for the vertices of the hypercube $H=\{-1,1\}^{d}\subset\mathbb{R}^{d}$, a Steiner tree is necessarily full. By Lemma 1, a Steiner tree $\mathcal{S}t$ for $H$ lies in the convex hull of $H$, and for any points $x\in H$, $y,z\in\operatorname{conv}H$ the angle $\angle yxz$ is non-obtuse. Hence $x$ cannot have degree 2 or 3 in $\mathcal{S}t$.
@@
$\mathcal{H}^{1}(\mathcal{S}t_{r})=\mathcal{H}^{1}(\mathcal{S}t_{r}\cap B_{1}(o))\geq\frac{2(2^{d}-1)}{\sqrt{3d}},$
where $\mathcal{S}t_{r}$ is a Steiner tree for a rescaled hypercube, inscribed in $B_{1}(o)$ (so the scale factor is $1/\sqrt{d}$).
Now let us show that a main part of this example belongs to a smaller ball $B_{\rho}(o)$, where $\rho$ is close to 1.
Lemma 7.
Let $\mathcal{S}t$ be a Steiner tree for $A=\left\{\frac{-1}{\sqrt{d}},\frac{1}{\sqrt{d}}\right\}^{d}$. Then
$H^{1}(\mathcal{S}t\cap B_{\rho}(o))\geq\frac{2(2^{d}-1)}{\sqrt{3d}}\cdot\left(1-2\sqrt{3d(1-\rho^{2})}\right),$
provided that $1\geq\rho\geq 1-\frac{1}{10d}$.
Proof.
Let $S$ be a connected component of the set $\left[\frac{-1}{\sqrt{d}},\frac{1}{\sqrt{d}}\right]^{d}\setminus B_{\rho}(o)$. Clearly, $S$ consists of points with the same signs of coordinates (so there are $2^{d}$ such components); indeed, any point with at least one zero coordinate inside the hypercube $\frac{-1}{\sqrt{d}},\frac{1}{\sqrt{d}}]^{d}$ has distance at most
$\sqrt{0+1/d+\dots+1/d}=\sqrt{\frac{d-1}{d}}<1-\frac{1}{10d}$
from $o$ and thus belongs to $B_{\rho}(o)$.
Without loss of generality, let $S$ contains point $v$ with all coordinates equal to $1/\sqrt{d}$. Then the distance from $v$ to any point from the set $\partial S\cap\partial B_{\rho}(o)$ is at most
$\sqrt{\left(\frac{1}{\sqrt{d}}-x_{1}\right)^{2}+\dots+\left(\frac{1}{\sqrt{d}}-x_{d}\right)^{2}}\leq\sqrt{1-x_{1}^{2}-x_{2}^{2}-\dots-x_{d}^{2}}=\sqrt{1-\rho^{2}},$
where we use $0\leq x_{i}\leq 1/\sqrt{d}$ and $(a-b)^{2}\leq a^{2}-b^{2}$ for $0<b<a$.
Now recall that $\mathcal{S}t$ is a full tree and thus it consists of $2\cdot 2^{d}-3$ line segments. Since a segment intersects a sphere in at most two points, we have
$(\mathcal{S}t\cap\partial B_{\rho}(o))\leq 4\cdot(2^{d}-1.5).$
Apply Lemma 5 with $X=\mathcal{S}t\setminus B_{\rho}(o)$ and $Y$ being the union of all segments connecting points $\mathcal{S}t\cap\partial B_{\rho}(o)$ with a nearest vertex of a hypercube. The previous bounds give
$\mathcal{H}^{1}(X)\leq\mathcal{H}^{1}(Y)\leq 4(2^{d}-1.5)\sqrt{1-\rho^{2}}<4(2^{d}-1)\sqrt{1-\rho^{2}}.$
Thus
$\mathcal{H}^{1}(\mathcal{S}t\cap B_{\rho}(o))=\mathcal{H}^{1}(\mathcal{S}t)-\mathcal{H}^{1}(X)\geq\frac{2(2^{d}-1)}{\sqrt{3d}}-4\cdot(2^{d}-1)\sqrt{1-\rho^{2}}\geq\frac{2(2^{d}-1)}{\sqrt{3d}}\left(1-2\sqrt{3d(1-\rho^{2})}\right).$
@@ 4.2 Upper bound
By Theorem 3 there exists a set $S$ connecting $\partial B_{r}(x)\cap\mathcal{S}t$ with length at most $12rt_{r}^{\frac{d-2}{d-1}}$, so by Lemma 5 for $X=B_{r}(x)\cap\mathcal{S}t$ and $Y=S$, we have
@@
$t_{r}\geq\frac{L_{r}^{\frac{d-1}{d-2}}}{144}$
for every $r\in[\rho,1)$. Lemma 6 implies
$L_{R}\geq L_{\rho}+\int_{\rho}^{R}t_{r}\,dr\geq L_{\rho}+\frac{1}{144}\int_{\rho}^{R}L_{r}^{\frac{d-1}{d-2}}dr$ (2)
@@
$f(z)=a+b\int_{\rho}^{z}f(y)^{\alpha}\,dy,\quad a=L_{\rho},\quad b=\frac{1}{144},\quad\alpha=1+\frac{1}{d-2}.$
@@
$L_{\rho}\leq\left(\frac{144(d-2)}{1-\rho}\right)^{d-2}.$
@@
$(q-\rho)t\leq\mathcal{H}^{1}(B_{q}(x)\cap\mathcal{S}t)\leq\left(\frac{144d}{1-q}\right)^{d-2},$
@@
$t\leq\frac{144d^{d-2}(d-1)^{d-1}}{(1-\rho)^{d-1}(d-2)^{d-2}}<\frac{1}{2}\frac{144d^{d-1}}{(1-\rho)^{d-1}}.$
Let $u\in[\rho,q]$ be a radius for which $(\partial B_{u}(x)\cap\mathcal{S}t)=t$. The Steiner set $\mathcal{S}t$ has no terminals inside $B_{u}(x)$, so the number of segments in the forest $\mathcal{S}t\cap B_{u}(x)$ is at most
$2t-3\leq\frac{144d^{d-1}}{(1-\rho)^{d-1}}.$
@@
$\mathcal{H}^{1}(\mathcal{S}t_{\varepsilon}\cap B_{\rho\varepsilon}(x))\leq\mathcal{H}^{1}(\mathcal{S}t\cap B_{\rho\varepsilon}(x))\leq\varepsilon\left(\frac{144d}{1-\rho}\right)^{d-2}.$
@@ 4.3 Planar case
Now consider $\mathcal{S}t_{\varepsilon}=\mathcal{S}t\cap\overline{B_{1-\varepsilon}(o)}$, where $o$ is the origin and $\varepsilon>0$. By Corollary 1 $\mathcal{S}t_{\varepsilon}$ is a finite tree; let $\{p_{1},\dots,p_{n}\}=\mathcal{S}t\cap\partial B_{1-\varepsilon}(o)$. Corollary 5 and rescaling give that $\mathcal{H}^{1}(\mathcal{S}t_{\varepsilon})$ is the sum of $(1-\varepsilon)\cos\alpha_{k}$, where $\alpha_{k}$ is the angle between the radius $op_{k}]$ of $B_{1-\varepsilon}(o)$ and $\mathcal{S}t_{\varepsilon}$ at the point $p_{k}$. Clearly, $\alpha_{k}$ belongs to $0,\pi/2)$, so every summand is positive.
@@
If $x\in\omega$ is an accumulation point of $\mathcal{A}$, then it is an accumulation point from both sides. Formally, for every $\varepsilon>0$ every connected component of $(\omega\cap B_{\varepsilon}(x)\setminus\{x\})$ has a nonempty intersection with $\mathcal{A}$.
@@
Remark 2.
We cannot avoid this epsilondeltamanship because there is no analog of Corollary 4 for infinite trees. Indeed, an infinite sum of numbers on the complex unit circle does not necessarily converge. Also, in the proof of Lemma 4 it is possible to replace the summation over edges by the summation over vertices only if the sum of vertices (as complex numbers) converges absolutely. Again, this is not the case for terminals located on the complex unit circle.
@@
The length of a full Steiner tree with at least three terminals on a unit circle is at least $\sqrt{3}$.
@@
Suppose the contrary. Then there is a full Steiner tree $\mathcal{S}t$ with the length at most $\sqrt{3}(1-\varepsilon)$, which means that a finite full tree $S_{\varepsilon}$ inside $B_{1-\varepsilon}(o)$ also has length at most $\sqrt{3}(1-\varepsilon)$. We may assume that $S_{\varepsilon}$ has at least 3 terminals, otherwise take smaller $\varepsilon$.
So $S_{\varepsilon}$ is a locally minimal tree and one can apply Lemma 2 (Melzak reduction). By the pigeon-hole principle there are at least two choices of $p_{1},p_{2}$ and $q$, where $p_{1}$ and $p_{2}$ are terminals adjacent to the branching point $q$. Since $S_{\varepsilon}$ is a tree exactly one of the arcs $\breve{p_{1}p_{2}}$ does not contain any other terminals. The interiors of these arcs for two choices are disjoint, so one of them has angular measure smaller than $4\pi/3$, and so the Melzak point $p$ lies outside the circle $\partial B_{1-\varepsilon}(o)$. Also, this arc $\breve{p_{1}p_{2}}$ belongs to the quadrilateral $pp_{1}qp_{2}$ and so Melzak reduction in this case produces a tree $S^{\prime}_{\varepsilon}$. Note that $S^{\prime}_{\varepsilon}\cap B_{1-\varepsilon}(o)$ is a full tree with the length at most $\sqrt{3}(1-\varepsilon)$, so we can repeat the reduction until we obtain a tree with 3 terminals $a_{1}$, $a_{2}$ and $a_{3}$. If the diameter of $\{a_{1},a_{2},a_{3}\}$ is at most $\sqrt{3}(1-\varepsilon)$ then they can be covered by an arc $R$ of the angular measure $2\pi/3$; this gives that one of the angles, say $\angle a_{1}a_{2}a_{3}$, is at least $2\pi/3$. This contradicts the existence of a full tree (in this case it is a regular tripod) with terminals $a_{1}$, $a_{2}$ and $a_{3}$. Finally, the length of any connected set containing $\{a_{1},a_{2},a_{3}\}$ is at least $\operatorname{diam}\{a_{1},a_{2},a_{3}\}$. ∎
@@ 5 Open questions
We want to understand the behavior of the upper bounds in Theorem 2 and Corollary 2 with respect to $d$ and $1-\rho$. There is a large gap between the mentioned upper bounds and the lower bounds in the spirit of Lemma 7. One of the obstacles is the lack of nontrivial examples in high dimensions, see the discussion in [5].
Also, note that the results by Paolini and Stepanov work in a proper metric space $\mathcal{X}$. The quantitative picture in various metric spaces is expected to be diverse, as we can see even in the planar and non-planar Euclidean cases.
Quantitative regularity of other one-dimensional problems is also of interest. One of the challenging examples is Branched Optimal Transport (Gilbert–Steiner problem). The Gilbert–Steiner problem is a generalization of the Steiner tree problem and specific optimal mass transportation, which allows the use of additional (branching) point in a transport plan. A specific feature of the problem is that the cost of transporting a mass $m$ along a segment of length $l$ is equal to $m^{p}\cdot l$ for a fixed $0\leq p<1$ and segments may end at points not belonging to the supports of given measures (branching points). Our arguments in Theorem 2 and Corollary 2 can be repeated word for word for the case $p=0$. For more details, see [1].
@@ Acknowledgments.
The authors are grateful to Ilya Bogdanov for his questions, which led to clarification of the statements.
Funding.
This work started at Summer Scientific Workshop ‘‘Introduction to Modern Mathematics and Related Scientific Projects’’ supported by the Bulgarian Ministry of Education and Science, Scientific Programme ‘‘Enhancing the Research Capacity in Mathematical Sciences (PIKOM)’’, No. D01-88/20.06.2025. The work of P. Prozorov was supported by the Russian Science Foundation (project no. 25-11-00058). Y. Teplitskaya is supported by the French National Research Agency (ANR) under grant ANR-21-CE40-0013-01 (project GeMfaceT).
Data availability statement.
This manuscript has no associated data.
Conflict of Interest.
The authors have no relevant financial or non-financial interests to disclose.